for the following function, c computes a value, where if you add millions of dollars to the value, the result is the cost of implementing a city recycling project when x, as a perentage(not its decimal equivalent) citizens participate. C(x)= 1.5x/100-x
a. Using this model, determine the cost if 60% of the citizen's participate.
b. Using this model, find the percentage of participation that can be expected if $5 million is spent on this recycling project? Set up an equation and solve algebraically. Round to the nearest whole percent.

Answer by stanbon(75887)

C(x)= 1.5x/100-x
a. Using this model, determine the cost if 60% of the citizen's participate.
C(60) = 1.5*60/(100-60) = 90/40 = 9/4 = $2.25 million
---------------------------------------------------------

b. Using this model, find the percentage of participation that can be expected if $5 million is spent on this recycling project?
5 = 1.5x/(100-x)
500 -5x = 1.5x
6.5x = 500
x = 76.92%
================
Cheers,
Stan H.



Set up an equation and solve algebraically. Round to the nearest whole percent.